DeWitt, Corp just paid a dividend of $1 and the dividend will be growing at a constant rate of 20% for 2 years, and after that it will be growing at 7%. What is the intrinsic value of DeWitt's stock if investors require a rate of return of 9.5%? (Do not round intermediate calculations. Round your answer to 2 decimal places.)
Answer #1

Dividend for next 2 years

D1= 1*120%=1.2

D2=1.2*120%= 1.44

Horizon value=P2= D3/(k-g) = 1.44*107%/(9.5%-7%) = 61.63

Value of stock = D1/(1+k)+ (D2+P2)/(1+k)^2

= 1.2/1.095^1+(1.44+61.63)/1.095^2

=53.7